A father's age is three times the sum of the ages of his two children. In four years, he will be twice as old as the sum of their ages. How old is the father? Only an algebraic solution will be accepted.
Answer by ankor@dixie-net.com(22740)

let f = father's present age
let s = sum of the children's ages now
:
A father's age is three times the sum of the ages of his two children.
f = 3s
In four years, he will be twice as old as the sum of their ages.
f + 4 = 2(s+8); both children will age 4 yrs
f = 4 = 2s + 16
f = 2s + 16 - 4
f = 2s + 12
Replace f with 3s from the 1st equation
3s = 2s + 12
3s - 2s = 12
s = 12 is the sum of his childen
then
f = 3(12)
f = 36 yrs is the father's present age